OpenStudy (anonymous):

Algebra 2 HELP! PLEASE!! Digital downloads of new music releases cost $3.00 per song. Digital downloads of older music cost $2.00 per song. Marissa buys a total of 12 songs and spends $32.00. Represent the problem as a matrix equation. Let x be the number of new songs and y be the number of older songs. What is the solution of the matrix equation?

OpenStudy (anonymous):

\[x+y = 12\] \[3x+2y=32\]\[\left[\begin{matrix}1 & 1 \\ 3 & 2\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}12 \\ 32\end{matrix}\right)\]